Minda invested her 5,000 earnings from working abroad in two banks, part at 8% annual interest and the rest at 9%. If the total annual interset for both dollar deposits is $435, how much was invested at each rate?
Answer by stanbon(75887)

Minda invested her 5,000 earnings from working abroad in two banks, part at 8% annual interest and the rest at 9%. If the total annual interset for both dollar deposits is $435, how much was invested at each rate?
-----------------
Equations:
e + n = 5000 dollars
8e+9n = 48500
-------------------------
Modify::
8e + 3n = 8*5000
8e + 9n = 48500
-----
Subtract and solve for "n"::
6n = 8500
n = $1416.67 (amt. invested at 9%)
Solve for "e":
e = 5000 - n = $3583.67
------------
Cheers,
Stan H.
-----------